OpenStudy (anonymous):

Compare 3^pi and pi^3

OpenStudy (anonymous):

What do you get when you compute it? \[\pi ^3 \implies (\pi)(\pi)(\pi)\] \[\huge3^{(\pi)} = ?\]

OpenStudy (kendricklamar2014):

\[3^\pi = 31.544281\] \[\pi^3 = 31.006277\]

OpenStudy (kendricklamar2014):

Comparison: They both equal 31

OpenStudy (anonymous):

sorry .! i need whether <, >

OpenStudy (anonymous):

So what do you think, Kendrick did the computation for you, that should be easy.

OpenStudy (kendricklamar2014):

> = Greater-Than Sign < = Less-Than Sign

OpenStudy (anonymous):

@iambatman it is not good enough for me.>!

OpenStudy (kendricklamar2014):

What seems to be the problem @Joseph91

OpenStudy (kendricklamar2014):

Which is bigger @Joseph91 : \[31.544281 ~ or ~ 31.006277\]

OpenStudy (anonymous):

the problem is like compare x^y and y^x when x,y>0 you should not use calculator

OpenStudy (triciaal):